The units’ digit of a two digit number is twice the tens’ digit. If the digits
of the number were reversed the resulting number would be six less than
twice the original number. Find the original number.
A. 12; B. 24; C. 36; D. 48; E. none

Answer by stanbon(48558)

The units’ digit of a two digit number is twice the tens’ digit. If the digits
of the number were reversed the resulting number would be six less than
twice the original number. Find the original number.
----
Let the original number be 10t+u
The reversed-digit number is 10u+t
---------------------
Equations:
u = 2t
10u+t = 2(10t+u)-6
--------------
Substitute for "u" and solve for "t":
10(2t)+t = 2(10t+2t)-6
21t = 24t-6
3t = 6
t = 2 (the ten's digit)
--------------------------------
Since u = 2t, u = 4 (the unit's digit)
======================================
Original number = 24
==========================================
Cheers,
Stan H.
A. 12; B. 24; C. 36; D. 48; E. none